# The Matthew Effect of AI Programming Assistants: A Hidden Bias in Software Evolution

### v2

@@ 6 Conclusion
This study provides the first large-scale empirical evidence of the Matthew effect in AI programming assistants, demonstrating how LLMs systematically amplify existing popularity hierarchies among programming languages and frameworks. Our findings reveal that mainstream technologies consistently achieve higher success rates in code generation, while niche and emerging alternatives face disproportionate failure rates that could potentially stifle innovation. This bias toward popular ecosystems may increasingly influence technical decision-making in industry, where AI compatibility could become as significant as traditional considerations like performance and scalability. Moving forward, we plan to expand our benchmarks into broader domains, investigate collaborative multi-agent development scenarios, and develop methods to counteract ecosystem homogenization through diversity-aware training and inference strategies.
@@
We ensure reproducibility by releasing the complete benchmark dataset, prompt templates, and evaluation code. Details of the benchmark composition are given in Appendix A, prompt and code extraction pipelines in Appendix B, and experimental infrastructure in Table 6. All generated solutions, evaluation results, and scripts are provided in the supplementary material.
@@
This study used large language models (LLMs) as a core tool within the experimental process to generate and evaluate code samples. Our experiments involved calling the APIs of multiple LLM models (including GPT-4o-mini, DeepSeek-V3, Gemini-2.0-Flash, Gemini-2.5-Flash, and Qwen3-Turbo), making a total of over 120,440 code generation requests. These models were used to generate solutions for thousands of algorithmic programming tasks and hundreds of framework selection tasks, systematically investigating how AI-assisted programming impacts the software ecosystem. We followed a strict VibeCoding protocol, where the LLMs acted as autonomous agents to produce runnable code through iterative feedback on error messages.
@@ A.1 LeetCode Benchmark Composition
The complete dataset, including all question IDs, difficulty levels, and topic tags, is available in the supplementary material file leetcode_benchmark.csv.
@@ A.2 Data Availability and Reproducibility
We provide complete supplementary materials including the LeetCode benchmark dataset, all AI-generated solutions (from five LLM models across eight languages), and LeetCode evaluation results. The experiment is fully reproducible using our provided codebase, though we note two practical constraints: (1) the evaluation requires 6-9 weeks due to rate limiting and scale (120,440 code generations), and (2) reproduction incurs significant API costs (\$1500-\$2,000 USD). All materials are structured for straightforward replication.
@@ D.2 Additional Model Results for Figure 1
Due to space constraints in the main body, the success rate (Pass@1) curves for two models (Gemini-2.0-Flash and Qwen3-Turbo) across all difficulty levels and eight programming languages are presented here in Figure A1. The trend observed in the main text—where performance degrades significantly for less popular languages (Erlang, Racket) especially on harder problems—is consistent across all five evaluated models.
@@ E.2 License for the Dataset of Authored Tasks
The benchmark tasks authored by the authors of this study are released under the MIT License to encourage future research. The license file is included in the supplementary material.

### v3

@@ 6 Conclusion
This study provides the first large-scale empirical evidence of the Matthew effect in AI programming assistants, demonstrating how LLMs systematically amplify existing popularity hierarchies among programming languages and frameworks. Our findings reveal that mainstream technologies consistently achieve higher success rates in code generation, while niche and emerging alternatives face disproportionate failure rates that could potentially stifle innovation. We emphasize that technical decision-making is multi-dimensional; AI compatibility is not a universal veto that overrides established factors like runtime performance. However, our results quantify a substantial ’AI Productivity Tax’ for niche languages. This creates a hidden friction consistent with Matthew-effect dynamics, which may disproportionately influence new projects and long-term ecosystem diversity. Moving forward, we plan to expand our benchmarks into broader domains, investigate collaborative multi-agent development scenarios, and develop methods to counteract ecosystem homogenization through diversity-aware training and inference strategies.
@@
We ensure reproducibility by releasing the complete benchmark dataset, prompt templates, and evaluation code. Details of the benchmark composition are given in Appendix A, prompt and code extraction pipelines in Appendix B, and experimental infrastructure in Table 6. All code, prompts, and released artifacts are available at our public repository: https://github.com/FrankGGu/The-Matthew-Effect-of-AI-Programming-Assistants. For large files, we use GitHub Releases under the same repository.
@@
This study used large language models (LLMs) as a core tool within the experimental process to generate and evaluate code samples. Our experiments involved calling the APIs of multiple LLM models (including GPT-4o-mini, DeepSeek-V3, Gemini-2.0-Flash, Gemini-2.5-Flash, and Qwen3-Turbo), making a total of over 135,495 code generation requests. These models were used to generate solutions for thousands of algorithmic programming tasks and hundreds of framework selection tasks, systematically investigating how AI-assisted programming impacts the software ecosystem. We followed a strict VibeCoding protocol, where the LLMs acted as autonomous agents to produce runnable code through iterative feedback on error messages.
@@
Ethics Statement
This study relies exclusively on data from LeetCode.cn, a platform independently operated by Lingkou Network (Shanghai) Co., Ltd. and governed by its own Terms of Service (https://leetcode.cn/terms-c). The intellectual-property attorney confirmed that retrieving publicly available problem metadata, without bypassing technical protection, while respecting the Robots protocol, and for non-commercial academic research, is compliant with both Chinese law and the LeetCode.cn Terms of Service. We additionally confirmed our access pattern and usage constraints with LeetCode.cn staff.
All accounts used for submitting model-generated solutions were manually created, used solely for distributing evaluation load, and operated strictly within normal rate limits. We did not attempt to obtain platform benefits, circumvent protections, or access any non-public content. We do not process, collect, or store personal data, user-generated content, or paid materials from LeetCode.cn. We do not redistribute proprietary problem statements or other restricted content. Researchers with enterprise access may further reproduce our evaluation pipeline through the official LeetCode API endpoint (https://leet.ai ), which imposes no strict rate limits and supports all languages required in our study.
@@ A.2 Data Availability and Reproducibility
We release the benchmark construction scripts, prompt templates, evaluation pipeline, and analysis code at: https://github.com/FrankGGu/The-Matthew-Effect-of-AI-Programming-Assistants. The experiment is fully reproducible using the provided codebase, though we note two practical constraints: (1) the evaluation requires 7–10 weeks due to rate limiting and scale (135,495 code generations), and (2) reproduction incurs significant API costs (\$1800-\$2,000 USD).
@@ D.2 Additional Model Results for Figure 1
Due to space constraints in the main body, the success rate (Pass@1) curves for two models (Gemini-2.0-Flash and Qwen3-Turbo) across all difficulty levels and eight programming languages are presented here in Figure 6. The trend observed in the main text—where performance degrades significantly for less popular languages (Erlang, Racket) especially on harder problems—is consistent across all five evaluated models.
@@ D.3 Experimental Screenshots
D.4 Potential Extension: Latent Factor Analysis of Task–Stack Outcomes
[…]
@@ E.2 License for the Dataset of Authored Tasks
The license is provided in our repository: https://github.com/FrankGGu/The-Matthew-Effect-of-AI-Programming-Assistants